# Data Augmentation using LLMs: Data Perspectives, Learning Paradigms and Challenges

### v1

@@ Appendix A Tasks
In this section, we classify the works based on the task that data augmentation is used to solve. We categorize the tasks as fundamental tasks, understanding based tasks, generation tasks, multi-modal tasks, and other tasks. We elaborate on the fundamental tasks in A.1, which includes text classification, machine translation, and sequence tagging. We explain the understanding and inference tasks in A.2, with question answering, logical reasoning, and natural language inference. We discuss the generation and interaction tasks in A.3, which includes summarization, data-to-text NLG, open-ended and conditional generation, and dialog. We introduce multi-modal tasks in A.4, which includes story telling and VQA tasks. Lastly, we summarize some other tasks that do not belong to above mentioned categories in A.5.
@@ A.1 Fundamental Tasks
For machine translation, data augmentation using LLMs serves two major purposes. The first purpose is enlarge data corpus to faciliate training, and the second purpose is to enhance evaluation quality. To achieve the first purpose, Oh et al. (2023) explore prompt-based data augmentation approaches that leveraging ChatGPT. It compares three paradigms "paraphrase", "multi-target", and "storytelling" and generates synthetic data on Korean-German language pairs from AI-hub dataset. To achieve the second purpose, DATScore Eddine et al. (2022) utilizes augmented data translated from the source and reference texts using ChatGPT. Experiments on WMT17 and WMT18 shows that DATScore correlates better with human judgements than BLEU, BERTScore Zhang et al. (2019), and BARTScore Yuan et al. (2021) on machine translation task.
@@ A.2 Understanding and Inference Tasks
  • counterfactual example: Li et al. (2023d) employ ChatGPT to generate counterfactual samples and counterfactual labels for each original sample. By merging the original data with the counterfactual data, an augmented dataset is formed to enhance the small models’ performance on the NLI task. Disco Chen et al. (2023d) identifies potential locations for performing counterfactual perturbations on the target instances, and then prompts GPT-3 to generate perturbations for NLI task.
@@ A.5 Other Tasks
One key area of data augmentation’s usage of tooling is in code generation and assistance. LLMs are employed to assist in writing and optimizing code, offering suggestions, debugging help, and even generating code snippets based on user requirements. Dong et al. (2023b) explore the existing data augmentation techniques from NLP and graph learning for enhancing training quality in source code learning. After reviewing and categorizing the literature, 11 data augmentation methods from NLP and graph learning are identified as potentially applicable to source code learning. Zhang et al. (2023d) enhance code generation by integrating tool usage information via an automated annotation method and API search tools, showing significant improvements in benchmarks. Despite its smaller size, ToolCoder’s performance is comparable to GPT-3.5, illustrating the benefits of embedding programming tools into code generation models.
@@ B.1 Clinical
Medical question answering is another field of interest. Guo et al. (2023) develop smaller efficient language models with augmentation data generated by LLMs on the PubMedQA dataset. The proposed method uses GPT-3.5 Turbo and GPT-4 to rewrite existing QA pairs or generate new pairs via zero-shot prompting. Results show that the small fine-tuned model with the data augmentation approach outperforms few-shot GPT4 on the PubMedQA dataset. For medical dialogue applications, GPT-3-ENS is introduced to create high-quality synthetic training data via low-shot learning and an ensemble method. Together with the human labeled data, the generated data help yield more performant models in terms of medical accuracy and coherency Chintagunta et al. (2021).

### v2

@@ 7 Limitations
In this section, we classify the works based on the task that data augmentation is used to solve. We categorize the tasks as fundamental tasks, understanding based tasks, generation tasks, multi-modal tasks, and other tasks. We elaborate on the fundamental tasks in section A.1, which includes text classification, machine translation, and sequence tagging. We explain the understanding and inference tasks in section A.2, with question answering, logical reasoning, and natural language inference. We discuss the generation and interaction tasks in section A.3, which includes summarization, data-to-text NLG, open-ended and conditional generation, and dialog. We introduce multi-modal tasks in section A.4, which includes story telling and VQA tasks. Lastly, we summarize some other tasks that do not belong to above mentioned categories in section A.5.
@@ A.1 Fundamental Tasks
For machine translation, data augmentation using LLMs serves two major purposes. The first purpose is enlarge data corpus to faciliate training, and the second purpose is to enhance evaluation quality. To achieve the first purpose, Oh et al. (2023) explore prompt-based data augmentation approaches that leveraging ChatGPT. It compares three paradigms ”paraphrase”, ”multi-target”, and ”storytelling”and generates synthetic data on Korean-German language pairs from AI-hub dataset. To achieve the second purpose, DATScore Eddine et al. (2022) utilizes augmented data translated from the source and reference texts using ChatGPT. Experiments on WMT17 and WMT18 shows that DATScore correlates better with human judgements than BLEU, BERTScore Zhang et al. (2019), and BARTScore Yuan et al. (2021) on machine translation task.
@@ A.2 Understanding and Inference Tasks
  • counterfactual example: Li et al. (2023d) employ ChatGPT to generate counterfactual samples and counterfactual labels for each original sample. By merging the original data with the counterfactual data, an augmented dataset is formed to enhance the small models’ performance on the NLI task. Disco Chen et al. (2023e) identifies potential locations for performing counterfactual perturbations on the target instances, and then prompts GPT-3 to generate perturbations for NLI task.
@@ A.5 Other Tasks
One key area of data augmentation’s usage of tooling is in code generation and assistance. LLMs are employed to assist in writing and optimizing code, offering suggestions, debugging help, and even generating code snippets based on user requirements. Dong et al. (2023c) explore the existing data augmentation techniques from NLP and graph learning for enhancing training quality in source code learning. After reviewing and categorizing the literature, 11 data augmentation methods from NLP and graph learning are identified as potentially applicable to source code learning. Zhang et al. (2023d) enhance code generation by integrating tool usage information via an automated annotation method and API search tools, showing significant improvements in benchmarks. Despite its smaller size, ToolCoder’s performance is comparable to GPT-3.5, illustrating the benefits of embedding programming tools into code generation models.
@@ B.1 Clinical
Medical question answering is another field of interest. Guo et al. (2023) develop smaller efficient language models with augmentation data generated by LLMs on the PubMedQA dataset. The proposed method uses GPT-3.5 Turbo and GPT-4 to rewrite existing QA pairs or generate new pairs via zero-shot prompting. Results show that the small fine-tuned model with the data augmentation approach outperforms few-shot GPT4 on the PubMedQA dataset. For medical dialogue applications, GPT-3-ENS is introduced to create high-quality synthetic training data via few-shot learning and an ensemble method. Together with the human labeled data, the generated data help yield more performant models in terms of medical accuracy and coherency Chintagunta et al. (2021).